After successful new product introduction Elaine decided to establish two muffin top mobile stands to create higher product exposure to potential customers in other areas. After some time, she conducted a market survey, which asked 300 randomly selected people about the frequency of their muffin top purchases from those mobile stands. Survey results are shown in the table below:

purchase frequency

mobile stand 1

(B1

mobile stand 2

B2

totals:
frequency A1 81 63 144
Rarely A2 48 42 90
Never A3 36 30 66
totals: 165 135 300

A. What is the probability of randomly selecting a person, who frequently buys muffin tops at the mobile stand 2?

B. What is the probability of randomly selecting a person, who never buys muffin tops at the mobile stand 1?
Answer #1

a) P(Frequently and Mobile stand 2) = 63/300 = 0.210

b) P(Never and Mobile stand 1) = 36/300 = 0.120